# SCAIL: Towards Studio-Grade Character Animation via In-Context Learning of 3D-Consistent Pose Representations

### v2

@@ Abstract
Achieving character animation that meets studio-grade production standards remains challenging despite recent progress. Existing approaches can transfer motion from a driving video to a reference image, but often fail to preserve structural fidelity and temporal consistency in wild scenarios involving complex motion and cross-identity animations. In this work, we present SCAIL (a framework toward Studio-grade Character Animation via In-context Learning), a framework designed to address these challenges from two key innovations. First, we propose a novel 3D pose representation, providing a more robust and flexible motion signal. Second, we introduce a full-context pose injection mechanism within a diffusion-transformer architecture, enabling effective spatio-temporal reasoning over full motion sequences. To align with studio-level requirements, we develop a curated data pipeline ensuring both diversity and quality, and establish a comprehensive benchmark for systematic evaluation. Experiments show that SCAIL achieves state-of-the-art performance and advances character animation toward studio-grade reliability and realism. Code and model are available at zai-org/SCAIL.
Figure 1: We propose SCAIL, a studio-grade character animation framework. SCAIL enables high-fidelity character animation under diverse and challenging conditions, including large motion variations, stylized characters, and multi-character interactions.
@@ 1 Introduction
High-fidelity character animation has tremendous potential for film production. Conventional filmmaking pipelines rely on a complicated workflow—involving motion capture, rigging, rendering—that demands expensive hardware and significant expert labor. Recently, the emergence of video generation models [12, 2, 47, 36] introduces a new paradigm: given a reference image and a driving video, such models can directly synthesize a photorealistic animation that follows the target motion in driving video, dramatically lowering the production barrier for high-quality character animation. Similar to real-world production pipelines, current methods [44, 15, 52, 49, 31, 30, 38, 4] typically begin by extracting skeletal motion sequences from the driving video as a form of “motion capture”, and then inject this information into a video generation model to perform “rigging” and “rendering”. However, in studio-grade production setting, these methods often struggle with challenging scenarios, such as complex motions (e.g., turning, rolling, flipping), multi-person interactions (e.g., dancing, hugging, fighting), and cross-domain animation where the reference and driving subjects differ significantly in appearance or body shape. As a result, generated videos frequently exhibit artifacts including distorted appearance, implausible body poses, and incorrect occlusions. These limitations reveal that current skeletal pose representations and generation models fail to adequately capture 3D structure of driving motions, inter-character spatial and occlusion relationships, and temporal correlations of motion sequences.
@@
To address the above limitations, we present SCAIL, a framework that revisits two core technical bottlenecks: (1) how to build a pose representation that effectively bridges driving and generated videos with unambiguous and accurate motion encoding, and (2) how to inject pose control in a way that enables the model to capture spatio-temporal motion structures. First, for motion representation, prior works [15, 49, 31, 38, 17, 4] typically rely on 2D skeletons (extracted from DWPose [46], ViTPose [43], etc.), which suffer from noisy predictions in complex motions and cannot encode occlusion. Some recent methods [52, 37] adopt SMPL-based [20] controls, which offer strong 3D human priors but have limited augmentation flexibility and cause identity leakage due to person-specific shape parameters. To address these issues, we propose a novel 3D pose representation. Specifically, we connect estimated 3D human keypoints according to skeletal topology and represent bones as spatial cylinders. The resulting 3D skeleton is rasterized to obtain 2D motion guidance signals. Building on this representation, we further design pose augmentation and retargeting strategies, enabling seamless scaling across diverse characters and scenarios. Second, regarding pose injection, a common approach in diffusion transformer (DiT) [23] model is to concatenate conditions channel-wise [51, 4]. However, this only provides local motion cues without capturing global motion dependencies. To overcome this, we propose a full-context pose injection mechanism within a DiT-based architecture. This design allows the model to attend to the entire pose sequence when generating each frame, enabling it to reason about motion context across time, and better capture high-level motion semantics. As a result, our method supports robust motion transfer across diverse body types and visual domains.
To achieve studio-level generation quality, we build a data curation pipeline ensuring both diversity and fidelity. Our pipeline integrates multi-domain sources covering various character types and motion styles. During curation, automatic filtering removes low-quality clips based on human presence and clarity, and motion-based metrics are further computed to identify motion-rich samples. A final manual stage further selects a finetuning set of superior quality.
@@
We also observe that current evaluations lack a comprehensive benchmark that adequately reflects production-level requirements. To address this gap, we propose Studio-Bench. It consists of two parts: the first evaluates motion adherence and structural integrity under complex single- and multi-person actions, while the second measures model generalization when the reference image and driving video differ in identity or domain. Our Studio-Bench covers challenging real-world cases and provides a realistic and rigorous measure of studio-level generation capability.
@@
Our main contributions can be summarized as follows: (1) We propose a scalable 3D pose representation that unifies the strengths of 2D skeleton and SMPL, serving as a robust motion-driving signal. (2) We inject driving-pose controls via in-context reasoning to enable effective spatiotemporal motion modeling, yielding superior results in complex and multi-person scenarios. (3) We construct a pipeline for curating high-quality, diverse training data, and establish a comprehensive Studio-Bench for systematic evaluation. (4) Our SCAIL framework achieves state-of-the-art performance over existing baselines, and advances character image animation toward production-level readiness.
@@ 2.1 Diffusion Models for Video Generation
2.2 Character Image Animation
@@ 3.1 Preliminaries
3D-Consistent Pose Representation. Recent methods [15, 30, 38, 17, 4] that employ 2D keypoint-based skeletons as motion controls perform well on simple actions. However, they often struggle under complex studio-level production scenarios, leading to abnormal human kinematics and violation of motion physics. 2D keypoints cannot encode 3D occlusion, making it difficult for models to generate realistic motion videos from incomplete or ambiguous pose signals. Another challenge arises from the discrepancy between the reference image and driving pose, which often differ in body shape or identity. Previous attempts [30, 38] mitigate the discrepancy through 2D skeleton scaling in training or heuristic retargeting in testing, yet 2D-derived adaptation inherently suffers from deformation due to inconsistency with 3D motion dynamics and further amplifies estimation noise and inaccuracies.
We argue that an ideal driving pose representation should perceive depth and occlusion, remain identity-agnostic, and retain the potential to scale to multi-person settings. To achieve that, we employ NLFPose [26] to estimate 3D body keypoints $\mathcal{J}=\{\text{J}_{i,j}\}_{i=1,j=1}^{T,N}$, where $\tilde{\text{J}}_{i,j}$ denotes the predicted 3D joint $j$ at frame $i$, and connect them according to the skeletal topology in 3D space. We discard SMPL mesh-like representations, which are inherently human-specific and offer limited controllability for augmentation or retargeting, often resulting in identity leakage. To retain occlusion and depth cues lost in 2D projection, we render 3D skeletons as cylindrical segments to provide motion-driving signals.
To address the disparity between the reference image and the driving pose, the framework should enable precise, identity-agnostic motion control and transfer motion naturally respecting the initial conditions present in the reference image. To achieve these objectives, we introduce a two-stage optimization strategy leveraging the 3D prior of our representation. During inference, we perform spatial alignment using a projection loss ($\mathcal{L}_{\text{proj}}$) to remap the predicted 3D driving poses to the reference 2D frame. This is to help the model anchor where the motion should be transferred to. Specifically, we optimize the following objective to refine the camera projection matrix $\tilde{\text{Cam}}^{3\times 3}$:
@@
where $\Pi(\cdot)$ is the perspective projection, $N$ is joint number, and $\text{P}^{\text{ref}}_{j}$ is the estimated 2D joints of the reference frame. During training, we decouple the augmentation process to body figure rescale and camera manipulation. To achieve identity decoupling, we synthesize a set of per-subject scale parameters $\text{S}_{i}$ and randomly sample the parameters during training to simulate different body proportions. Since the underlying estimated skeleton topology and relative bone directions remain consistent, the scaled skeletons still faithfully represent the same motions.
As the alignment stage in inference may introduces approximation error, we also exert augmentation on camera parameters ${\text{Cam}}^{3\times 3}$ to enhance the model’s robustness towards camera variance. Since in training we sample reference image from other parts of the video, we restrict the augmentation scale in camera manipulation within a limited range to provide anchor for where motion happens, so that our method can strike a balance between the controllability of location and natural motion transfer in further generation.
Full-Context Driving Pose Injection. Our model builds upon a DiT-based Image-to-Video (I2V) model [36] and inject pose control signals for motion guidance. The common practice for pose information injection is through channel concat [51, 4] or using a pose adapter [38], where the driving pose sequence is embedded and added to the noisy video latents. We first implement the channel concat approach as illustrated in Figure 4, following [51]. While this method shows decent pose-following capability in generation, we find it tend to generate unnatural human pose when the motion is complex despite conditioned on accurate and unambiguous pose. A typical example is the turning motion, where the model often fails to correctly distinguish between front and back views, generating awkward postures. We assume that this limitation stems from the per-frame addition scheme, which fails to provide sufficient temporal context for motion understanding. Since motion is inherently sequential, many actions can only be interpreted correctly within a temporal context. Therefore, we design a new strategy termed full-context pose injection, which enables the model to reason over the entire pose sequence. Specifically, we directly concatenates conditional pose tokens with noisy video tokens to facilitate spatial-temporal interactions between the two modality. As shown in Figure 4, this strategy handles motion correctly in complex scenarios.
@@
Shifted RoPE for Pose Context. Conventional pipelines typically extract poses that are spatially aligned with the original video. However, in our setting, augmentations such as scaling and camera transformations often prevent the extracted pose context to be spatially aligned with the driving video. To address this misalignment, we introduce Pose-Shifted RoPE mechanism, which enables the model to effectively retrieve driving signals from the augmented full-sequence pose representation. As illustrated in Figure 3, conventional 3D RoPE encodes position in the form $(t,h,w)$, where $t$ is the temporal frame index and $h,w$ denote spatial dimensions, each starting from zero. In our design, Pose-Shifted RoPE applies a shift along the width dimension specifically to pose representations. The positional vectors for each driving pose token $d$ are defined as:
@@
3.3 Data Collection
Data Source. Our dataset is composed of three primary data sources: (1) samples retrieved from our internal base model training data and other downstream tasks, (2) a large collection of high-resolution dance videos from Bilibili and YouTube, and (3) additional sports videos such as gymnastics and figure skating. To ensure diversity, we maintain a certain proportion of stylized content, including 3D and 2D animations from source (1), as well as MMD and Live2D animations from source (2).
Filtering and Pose Extraction Pipeline. We develop a curated pipeline to construct a high-quality dataset for character animation. To meet the requirements for generating studio-grade character animation, we have established the following criteria for data filtering: (1) the character should be the primary focus of each frame, (2) the person should exhibit explicit motion, and (3) the pose should be visually complete, covering either the upper body or the full body. We first detect human presence using YOLO [33] and discard clips where characters are not the main subjects. Next, 2D keypoints are extracted using DWPose [46], retaining only half-body or full-body videos. After this stage, multi-person videos can be further separated based on the number of bounding boxes detected by YOLO [33]. We further employ a large language model [7] to analyze captions, complementing the complex rule-based logic of multi-person motion bounding box checking. For multi-person videos, we employ Samurai [45] to track and split the mask of each character, generating multiple single-human video splits. Subsequently, 3D keypoints are extracted using NLFPose [26] for each video that contain one main character. To meet the motion criteria, we compute the motion speed from the estimated 3D keypoints and discard samples with limited motion, resulting in a clean and motion-rich training set. The motion speed $v$ is calculated as:
@@ 3.3 Data Collection
Finally, The 3D skeletons are represented as cylindrical segments in 3D space and rendered using the Taichi [16] renderer to preserve spatial relationships. For multi-person data, each split is independently processed by NLFPose to estimate 3D skeletons, which are then composed together in 3D space and rasterized on the 2D canvas. This procedure preserves inter-person occlusion relationships, benefiting from NLFPose’s accurate depth estimation. In practice, we find that our proposed multi-stage pipeline provides more accurate estimation than direct multi-person pose estimation methods like PromptHMR [39], especially in wild cases involving complex interaction.
@@ 4.2 Quantitative Evaluation
Figure 9: Visualization of ablation study.
@@ 4.4 Ablation Study
We conduct ablation studies on SCAIL-1.3B model to evaluate the contribution of each component. For reference, we also compare against the U-Net based MimicMotion [49], highlighting the advantage of our DiT-based architecture. In the main paper we mainly focus on ablations on model architecture, and more ablations on the pose condition would be provided in Appendix A2.
Ablation on Full-Context Driving Pose Injection. As shown in Table 2, replacing channel concatenation with the proposed full-context injection consistently improves all the metrics. We observe that generated videos exhibit fewer artifacts like limb tearing and structural collapse (shown in Figure 9), indicating that our context-aware injection strategy enables the model to capture global motion dependencies and reason plausible human poses. Our architecture is designed to enable extensive interaction between noisy video tokens and pose tokens, and experiments show the robustness this design achieves when facing pose-estimation errors. Specifically, case (2) shows that through this strategy, when the estimator misidentifies the left part and the right part and extracts a posture like a forward lunge in a running driving video, our model with full-context pose injection can still generate correct running posture based on the global motion, as conditions from the reference frame and estimations in other frames clearly reflect the semantic context of running along the river.
@@
SCAIL-1.3B w/o full-context 17.69 0.626 0.262 263.63
@@
Table 2: Ablation study conducted on SCAIL-1.3B. The w/o full-context denotes using the channel-concatenation strategy.

### v3

@@ Abstract
Achieving controllable character animation that meets studio-grade standards remains challenging despite recent progress. Existing approaches can transfer motion from a driving video to a reference image, but often fail to preserve structural fidelity and temporal consistency in wild scenarios involving complex motion and cross-identity animations. In this work, we present SCAIL (a framework toward Studio-grade Character Animation via In-context Learning), which is designed to address these challenges from two key innovations. First, we propose a novel 3D pose representation, providing a robust and flexible motion signal. Second, we introduce a full-context pose injection mechanism within a diffusion-transformer, enabling effective spatio-temporal reasoning over full motion sequences. To align with studio-grade requirements, we develop a curated data pipeline ensuring both diversity and quality, and establish a comprehensive benchmark for systematic evaluation. Experiments show that SCAIL achieves state-of-the-art performance and advances character animation toward studio-grade controlling. Code and model are available at zai-org/SCAIL.
Figure 1: We propose SCAIL, a character animation framework that enables high-fidelity character animation under diverse and challenging conditions, including large motion variations, stylized characters, and multi-character interactions.
@@ 1 Introduction
High-fidelity character animation has tremendous potential for film production. Conventional filmmaking pipelines rely on a complicated workflow—involving motion capture, rigging, rendering—that demands expensive hardware and significant expert labor. Recently, the emergence of video generation models [12, 2, 47, 36] introduces a new paradigm: given a reference image and a driving video, such models can directly synthesize animations that follow the target motion in driving video. Similar to real-world production pipelines, existing methods [44, 15, 52, 49, 31, 30, 38, 4] typically begin by extracting skeletal motion sequences from the driving video as a form of “motion capture”, and then inject this information into a video generation model to perform “rigging” and “rendering”. However, these methods often struggle with challenging scenarios, such as complex motions (e.g., turning, rolling, flipping), multi-person interactions (e.g., dancing, hugging, fighting), and cross-domain animation where the reference and driving subjects differ significantly in appearance or body shape, exhibiting distorted appearance or implausible limb occlusions. In this work, we identify these challenges as the primary bottleneck to achieving studio-grade controllable animation. These limitations reveal that current skeletal pose representations and pose-driving generation methods fail to adequately capture 3D structure of driving motions, inter-character spatial and occlusion relationships, and temporal correlations of motion sequences.
First, for motion representation, prior works [15, 49, 31, 38, 17, 4] typically rely on 2D keypoints (extracted from DWPose [46], ViTPose [43], etc.), which suffer from noisy predictions and cannot encode occlusion. Some works [52, 37] adopt SMPL [20] mesh controls, which offer strong 3D human priors but cause severe identity leakage. Second, regarding pose injection, a common approach in diffusion transformer (DiT) [23] model is to concatenate conditions channel-wise [51, 4]. However, we find that this method provides more local motion cues than capturing global motion dependencies.
@@
To address the above limitations, we present SCAIL, a framework that revisits two core technical bottlenecks: (1) how to build a pose representation that effectively bridges driving and generated videos with unambiguous and accurate motion encoding, and (2) how to inject pose control in a way that enables the model to capture spatio-temporal motion structures. We propose a novel 3D pose representation that rasterizes bones as spatial cylinders into the pixel plane, such representation can be augmented and retargeted to enable seamless controlling across diverse characters and scenarios. To prevent model from following local motion cues as a shortcut, we propose a full-context pose injection mechanism within a DiT [23]-based architecture. This design allows the model to attend to the entire pose sequence when generating each frame, enabling it to reason about motion context across time, and better capture high-level motion semantics.
@@
To obtain reliable video-pose pairs, we apply a segmentation-and-extraction approach to get 3D keypoints in human interactions. Building on that, a data curation pipeline ensuring both character diversity and motion complexity is established, with automatic filtering based on human presence and motion-based metrics and a final manual stage to select a finetuning set of superior quality. We also observe that current evaluations lack a comprehensive benchmark that adequately reflects production-level requirements. To address this gap, we propose Studio-Bench. It consists of two parts: the first evaluates motion adherence and structural integrity under complex single- and multi-person actions, while the second measures model generalization when the reference image and driving video differ in identity or domain. Our Studio-Bench covers challenging real-world cases and provides a realistic and rigorous measure of studio-level generation capability.
@@
Our main contributions can be summarized as follows: (1) We propose an identity-agnostic 3D pose representation that serves as motion-driving signal suitable for complex motions and multi-human interactions. (2) We inject driving-pose controls via in-context reasoning to enable effective spatiotemporal motion modeling, yielding superior results in complex and multi-person scenarios. (3) We construct a pipeline for curating high-quality, diverse training data, and establish a comprehensive Studio-Bench for systematic evaluation. (4) Our SCAIL framework achieves state-of-the-art performance over existing baselines, and advances character image animation toward production-level readiness.
@@ 2.1 Diffusion Models for Video Generation
2.2 Controllable Character Image Animation
@@ 3.1 Preliminaries
3D-Consistent Pose Representation. Recent methods [15, 30, 38, 17, 4] that employ 2D keypoint-based skeletons as motion controls perform well on simple actions. However, they often struggle under complex studio-level production scenarios, leading to abnormal human kinematics. As there exists discrepancy between the reference image and the driving pose (e.g. body shape variations), previous 2D-based methods [30, 38] mitigate such gap through 2D skeleton scaling during training or heuristic retargeting at test time. However, such adaptation inherently suffers from deformation due to inconsistency with 3D motion dynamics, further amplifying estimation noise and inaccuracies.
We design our 3D pose representation to retain depth and occlusion, deliver accurate motion and remain identity-agnostic. To achieve that, we employ NLFPose [26] to estimate 3D body keypoints $\mathcal{J}=\{\text{J}_{i,j}\}_{i=1,j=1}^{T,N}$, where $\tilde{\text{J}}_{i,j}$ denotes the predicted 3D joint $j$ at frame $i$, and connect them according to the skeletal topology in 3D space. We discard SMPL mesh-like representations, which are inherently human-specific and offer limited controllability for augmentation or retargeting, and render 3D skeletons as cylindrical segments to provide driving signals. We distinguish person identities by directly rendering skeletons with different hues instead of relying on hardcoded network designs [3].
To address the disparity between the reference image and the driving pose, we introduce a two-stage optimization strategy leveraging the 3D prior of our representation. Unlike previous methods [30] that applies 2D scaling in training augmentation and non-robust skeleton scaling rules in inference retargeting, we design training pose to be fully identity agnostic with accurate bone angles, and in inference we avoid altering the original estimated skeleton and only optimize the camera instead. To achieve identity decoupling during training, we synthesize a set of scale parameters $\mathcal{S}$, and randomly sample $S_{i}\in\mathcal{S}$ during training, each $S_{i}$ denoting a set of scaling parameter of each bones, simulating binding motion to characters with diverse body shapes. Since the underlying estimated skeleton topology and relative bone directions remain consistent, the scaled skeletons is temporally coherent and still faithfully represent the same motions.
In inference, we perform spatial alignment using a projection loss ($\mathcal{L}_{\text{proj}}$) to remap the predicted 3D driving poses to the reference 2D frame to help the model anchor where the motion should be transferred to, and optimize the training process to Specifically, we optimize the following objective to refine the camera projection matrix $\tilde{\text{Cam}}^{3\times 3}$:
@@
where $\Pi(\cdot)$ is the perspective projection, $N$ is joint number, and $\text{P}^{\text{ref}}_{j}$ is the estimated 2D joints of the reference frame.
As the alignment stage in inference may introduces approximation error, we simulates such error in training by directly applying such aligning method using the reference image and the first frame of the driving sequence (the reference image and driving sequence are sampled from different parts of the video) and exerting modest disturbance on camera parameters ${\text{Cam}}^{3\times 3}$ to enhance the model’s robustness towards camera variance. Such design strikes a balance between the controllability of location and robust motion transfer.
Full-Context Driving Pose Injection. Our model builds upon a DiT-based Image-to-Video (I2V) model [36] and inject pose control signals for motion guidance. The common practice for pose information injection is through channel concat [51, 4] or using a pose adapter [38], where the driving pose sequence is embedded and added to the noisy video latents. We first implement the channel concat approach as illustrated in Figure 4, following [51]. While this method shows decent pose-following capability in generation, we find it tend to generate unnatural human pose when the motion is complex despite conditioned on accurate and unambiguous pose. A typical example is the turning motion, where the model often fails to correctly distinguish between front and back views, generating awkward postures. We assume that this limitation stems from the per-frame addition scheme, which fails to provide sufficient temporal context for motion understanding. Since motion is inherently sequential, many actions can only be interpreted correctly within a temporal context. Therefore, we design full-context pose injection, which enables the model to reason over the entire pose sequence. Specifically, we directly concatenates conditional pose tokens with noisy video tokens to facilitate spatial-temporal interactions between the two modality. As shown in Figure 4, this strategy handles motion correctly in complex scenarios.
@@
Shifted RoPE for Pose Context. Conventional pipelines typically extract poses that are spatially aligned with the original video. However, in our setting, augmentations such as scaling and camera transformations often prevent the extracted pose context to be spatially aligned with the driving video. To address this misalignment, we introduce Pose-Shifted RoPE mechanism, which enables the model to effectively retrieve driving signals from the augmented full-sequence pose representation.
As illustrated in Figure 3, conventional 3D RoPE encodes position in the form $(t,h,w)$, where $t$ is the temporal frame index and $h,w$ denote spatial dimensions, each starting from zero. In our design, Pose-Shifted RoPE applies a shift along the width dimension specifically to pose representations. The positional vectors for each driving pose token $d$ are defined as:
@@
3.3 Data Pipeline
Pose Extraction. We adopt  [26] for 3D keypoints extraction. However, the default extraction pipeline often fails to detect correct limbs under occlusion, especially in multi-human interactions. To address so, we apply a segmentation-and-extraction approach to obtain reliable poses in human interactions. To be more specific, we employ Samurai [45] to track and split the mask of each character, generating multiple single-human video splits. Subsequently, 3D keypoints are extracted for each video that contain one main character. Finally, The 3D skeletons are represented as cylindrical segments in 3D space. Those cylindrical segments are then composed together in 3D space and rasterized on the 2D canvas. This procedure preserves inter-person occlusion relationships, benefiting from NLFPose’s accurate depth estimation. In practice, we find that our proposed multi-stage pipeline provides more accurate estimation than direct multi-person pose estimation methods like PromptHMR [39], especially in wild cases involving complex interaction. We follow the same extraction pipeline in inference.
Data Curation. To meet the requirements for generating studio-grade character animation, we have established the following criteria for data filtering: (1) the character should be the primary focus of each frame, (2) the person should exhibit explicit motion, and (3) the pose should be visually complete, covering either the upper body or the full body. We first detect human presence using YOLO [33] and discard clips where characters are not the main subjects. Next, 2D keypoints are extracted using DWPose [46], retaining only half-body or full-body videos. After this stage, multi-person videos can be further separated based on the number of bounding boxes detected by YOLO [33]. We further employ a large language model [7] to analyze captions, complementing the complex rule-based logic of multi-person motion bounding box checking. To compel the model to learn from complex motion dependencies, we compute the motion speed from the estimated 3D keypoints and discard samples with limited motion, resulting in a clean and motion-rich training set. The motion speed $v$ is calculated as:
@@ 4.2 Quantitative Evaluation
Figure 9: Visualization of ablation study. The w/o full-context denotes using the channel-concatenation strategy.
@@ 4.4 Ablation Study
We conduct ablation studies on SCAIL-1.3B model to evaluate the contribution of each component. For reference, we also compare against the U-Net based MimicMotion [49], highlighting the advantage of our DiT-based architecture. More ablations would be provided in Suppl. A2.
Ablation on the Pose Representation. To validate the effectiveness of our 3D-consistent pose representation, we compare with a 2D keypoints-based pipeline (extracted using DWPose [46] and denoted as w/ 2D Pose). Implementation details of the 2D keypoints-based baseline and qualitative comparisions are also provided in Suppl.A2. Results shows the pose representation is critical for the performance gain, and such improvements are more pronounced under full-context pose injection than under channel-concat scheme.
Ablation on Full-Context Driving Pose Injection. As shown in Table 2, replacing channel concatenation with the proposed full-context injection consistently improves all the metrics when providing the model with temporally coherent and information-rich conditions. Qualitatively, We observe that generated videos exhibit fewer artifacts like limb tearing and structural collapse (shown in Figure 9), indicating that our context-aware injection strategy enables the model to capture global motion dependencies and reason plausible human poses. Specifically, case (2) in  9 shows that through this strategy, when the estimator misidentifies the left part and the right part and extracts a posture like a forward lunge in a running driving video, our model with full-context pose injection can still generate correct running posture based on the global motion, as conditions from the reference frame and estimations in other frames clearly reflect the semantic context of running along the river.
@@
Channel Concat-1.3B 17.69 0.626 0.262 263.63
Channel Concat-1.3B w/ 2D Pose 17.12 0.624 0.282 296.23
@@
Table 2: Ablation study conducted on SCAIL-1.3B.